archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Details A first beta version of the incidents card is online http www ovhtelecom fr xdsl incident also accessible in the pad services status of the assistance section There s a French card which indicates per region the automatically detected incidents by our internal monitoring tools These data are accessible via the API xDSL To remind you when your line is impacted by an incident the customer interface displays already

    Original URL path: http://status.ovh.es/?do=details&id=6502 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Centre East Collection Closed 2013 05 30 4788 ADSL SDSL Modernization Backend Core FS 8719 DSLAM OVH Paris Closed 2013 05 30 4786 ADSL SDSL Modernization Backend Core FS 8717 Collection 13 Closed 2013 05 29 4779 ADSL SDSL Maintenance Backend Core FS 8710 DSLAM OVH Marseilles Closed 2013 05 29 4770 ADSL SDSL Maintenance Backend Core FS 8701 Eligibility Closed 2013 05 28 4755 ADSL SDSL Maintenance Backend Core FS 8686 NRA DRA13 Closed 2013 05 23 4605 ADSL SDSL Modernization Backend Core FS 8536 Bordeaux optic loop Closed 2013 04 25 4586 ADSL SDSL Modernization Backend Core FS 8519 lil 3 4 n5 Closed 2013 04 23 4566 ADSL SDSL Incident Backend Core FS 8497 LNS 1 PAR Closed 2013 04 19 4536 ADSL SDSL Incident Backend Core FS 8469 Collect 59 Closed 2013 04 17 4511 ADSL SDSL Maintenance Backend Core FS 8443 LNS 1 PAR Closed 2013 04 15 4497 ADSL SDSL Modernization Backend Core FS 8431 TG788vn modems Closed 2013 04 12 4496 ADSL SDSL Maintenance Backend Core FS 8430 Optical Loop Lyon Closed 2013 04 12 4492 ADSL SDSL Maintenance Backend Core FS 8426 Petite Couronne Paris optical loop Closed 2013 04 11 4396 ADSL SDSL Incident Backend Core FS 8330 LNS Paris Closed 2013 04 11 4477 ADSL SDSL Incident Backend Core FS 8411 Collect in the 35 Closed 2013 04 10 4471 ADSL SDSL Maintenance Backend Core FS 8405 Marseilles Optical Buckles Closed 2013 04 10 4433 ADSL SDSL Maintenance Backend Core FS 8367 Optic maintenance NRA Lille CAN59 and HAU59 Closed 2013 04 03 4400 ADSL SDSL Maintenance Backend Core FS 8052 Collection in 47 Closed 2013 03 28 4385 ADSL SDSL Modernization Backend Core FS 8319 DSLAM OVH Bordeaux Closed 2013 03 27 4360 ADSL SDSL Maintenance Backend Core FS

    Original URL path: http://status.ovh.es/?pagenum=5&string=&type=&sev=&dev=&due=&cat=&status=&order2=lastedit&sort=asc&sort2=desc&date=&project=&order=proj (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We re going to reset the PPP sessions ppp and shut down the lab router that does the MLPPP tonight at 0 00 Customers affected will switch to their own LNS Date Wednesday 05 March 2014 00 33AM Reason for closing Done Comments 3 Comment by OVH Wednesday 05 March 2014 00 32AM collecting subscribers sessions Comment by OVH Wednesday 05 March 2014 00 32AM clearing subscribers

    Original URL path: http://status.ovh.es/?do=details&id=6457 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Core Status CLOSED Percent Complete Details An optical maintenance is processing in Lyon by the fiber operator The following sites aren t secured BRO69 MEY69 DEC69 MPL69 SPR69 VEN69 BG469 SGL69 OUL69 PAR69 MON69 TOK69 LAL69 VIL69 MNE69 The site VEV69 is isolated The recovery is expected at 06 00am Date Friday 28 February 2014 14 59PM Reason for closing Done Comments 1 Comment by OVH Friday 28 February 2014

    Original URL path: http://status.ovh.es/?do=details&id=6442 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    feeder cards on ADVA chasssis of Telehouse 2 and Globalswitch No impact on the network is expected Date Wednesday 26 February 2014 12 31PM Reason for closing Done Comments 4 Comment by OVH Tuesday 25 February 2014 17 15PM We re starting the electrical intervention électrique on the Telehouse 2 site Comment by OVH Tuesday 25 February 2014 17 50PM The ADVA maintenance on the Telehouse site is finished Comment

    Original URL path: http://status.ovh.es/?do=details&id=6401 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details There has been an incident on the SFR equipment in Département 81 A ticket has been opened Date Tuesday 25 February 2014 15 37PM Reason for closing Done Comments 2 Comment by OVH Tuesday 25 February 2014 10 25AM The collect incident in Dept 81 FS 6413 is related to the same issue An update on the operator s core network router went wrong The

    Original URL path: http://status.ovh.es/?do=details&id=6414 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Status CLOSED Percent Complete Details There has been an incident on the SFR equipment in Département 12 A ticket has been opened Date Tuesday 25 February 2014 15 36PM Reason for closing Done Comments 2 Comment by OVH Tuesday 25 February 2014 10 25AM The collect incident in Dept 81 FS 6414 is related to the same issue An update on the operator s core network router went wrong The

    Original URL path: http://status.ovh.es/?do=details&id=6413 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    on the line tested Connection of a LiveBox with Orange ID directly on the strip Change of the FT LAC Change of the FT LNS After various tests were carried out we detected that a packet loss issue was still present behind each action The local loop and the OVH device are not causing the issue The issue lies with the Orange connection from start to finish The BAS was replaced by a BAS of the same model A replacement with a different module is scheduled We are calling Orange today to confirm an intervention date for changing the BAS and to plan follow up operations Comment by OVH Tuesday 06 May 2014 04 58AM Last week Orange has managed changing BAS but no effect was noticed on the packet loss In agreement with Orange we opened two incident tickets for two different customers with packet loss Orange planned interventions On the other hand following the request of Orange we have established an access list with and without packet loss This list contains information such as number of BAS connecting Our Orange speaker is analyzing and testing the access A reminder was made that day Finally last Monday the National Paris door was opened It is functional and now used for back up but will have no impact on the packet loss issue as the LNS is cleared of blame Comment by OVH Tuesday 06 May 2014 13 11PM We have been contacted by the Orange collect network service about working together After various tests the Orange engineer was unable to reproduce the bug on the Orange network We have arranged another meeting on site with a client so as to use their line undergoing packet loss to analyse and study the access Comment by OVH Wednesday 07 May 2014 13 49PM Orange has intervened on our client in order to carry out more tests At each intervention the tests carried out compliment the previous ones and new representaives are added to the loop At the same time Orange has carried out ATM and GE Giga Ethernet FTP download lab tests In these two cases gel during download packet loss was detected Orange has indicated that the problem may be located on their BAS or the OVH download servers which OVH contests On the other hand Orange has also performed ping tests and reports that a response is sometimes absent However Orange states that the machine ping isn t a priority and thus doesn t apply to this test Future actions listed by Orange are as follows OVH must provide an exhaustive list of customers with packet loss already done Orange is looking for a common point for domain names of this list Orange will mount a download test server so as to avoid using the OVH server Orange is putting its QoS teams in the loop Orange will make a new more realistic model in the lab and will relaunch the tests The malfunction is complex to analyse which explains the many interventions Comment by OVH Wednesday 21 May 2014 07 42AM Here is the result of telephone point lead today with Orange about packet loss correction Orange has reminded us of the results of tests that show the packet loss on our FTP server an FTP server which is independent but not on Oleane FTP server Orange doubts on FTP servers used Each of us must take the following actions to be discussed Wednesday night Orange should report to us the results of the analysis of the list of ND transmitted Orange must run tests from the client connection OVH will run tests on the Oleane FTP server to confirm or not their diagnosis In parallel with BAS a new lead has been identified the faulty routers network contacts of Orange and OVH will join the issue tomorrow Comment by OVH Thursday 22 May 2014 09 57AM Following the telephone review this evening the next actions will be as follows 1 Orange will temporarily lower the synchronisation of 0x xx xx xx xx which is undergoing packet loss on the Oleane FTP server 2 OVH will identify a customer with an ND undergoing packet loss and one without other than FT if possible and seek their permission to carry out new tests 3 OVH will suggest a set of tests to Orange in order to obtain a version approved by both parties by the next conf call 4 OVH and Orange will define a list of modems for carrying out these tests This list will be ready by the next conf call 5 Orange will suspend the investigations on the list of NDs provided by OVH in order to focus on these new actions The objective is agreeing on which tests are indisputable and ensuring that future movements on site will be managed by a representative that is following things at Orange We can t implicate customers and have them travelling for cancelled meetings or incomplete tests The next review will take place on Friday morning at 10 30 Comment by OVH Monday 26 May 2014 08 29AM Here is the report of the meeting with Orange about the packet loss 1 Synchronization falling of Orange had no impact 2 OVH has identified a customer who responds to all conditions namely an ND with packet loss and one without onthe same DSLAM 3 OVH and Orange have agreed on a test to check if a line is subject of the problem Generating 30Mbps UDP to the subscriber and check if the connection remains stable there may be loss of packets but not a full cut off 4 The list of compatible modems is no longer valid for the moment After this test day in collaboration with Orange on a vacuum line subject to packet loss here is the result The first step was to isolate precisely the conditions that lead to the loss of connection We noticed that when the line is heavily congested eg launch of TCP

    Original URL path: http://status.ovh.es/?do=details&id=6389 (2015-10-15)
    Open archived version from archive



  •